A charge q1=3.00 micro coulomb is at (0,0) and q2= -"5\/3" q1 is at 20cm mark on the positive x-axis. we also have a charge q3=8/5 q2 located at some unknown point on the x-axis. Where is q3 if the net force on q1 is 7.00N pushing q1 to the left?


1
Expert's answer
2021-08-09T16:24:24-0400

"F_2=\frac{kq_1q_3}{x^2},"

"F_1=\frac{kq_1q_2}{d^2},"

"F_2-F_1=F,"

"x=-\sqrt{\frac{kd^2q_1q_3}{Fd^2+kq_1q_2}}=-14.4~\text{cm}."